# Yi-Fang Wang, Chia-Ming Lee, Yi-Shiuan Chou

### v2

@@ 2.3 Proposed GCN with Graph Laplacian
Graph Convolutional Networks (GCNs) have emerged as a powerful tool for processing graph-structured data [47, 48], making them a suitable choice for handling the graph embedding obtained through feature entanglement. However, the noise level in each node can vary depending on the degree of distortion. For highly distorted or noisy nodes (i.e., those with entangled features primarily contributed by invalid facial images), it is beneficial to eliminate them to ensure stable and excellent performance in DeepFake video detection. To mitigate the impact of features from invalid facial images, we judiciously integrate the Graph Laplacian Smoothing Prior (GLSP), a well-established concept in Graph Signal Processing (GSP) [49], as a regularizer into the GCN to filter out highly noisy nodes [50]. It is important to note that GSP and GCN are distinct domains, with GSP focusing on the analysis and processing of signals defined on graphs, while GCN aims to learn representations by exploiting the graph structure. In this study, we ingeniously leverage the properties of GLSP from GSP and seamlessly incorporate it into the GCN framework, enabling end-to-end training without explicitly computing the eigendecomposition of the Graph Laplacian matrix.
Let $\mathbf{X}\in\mathbb{R}^{d\times c}$ denote the SFE feature matrix, where $d=Nwh$ represents the spatiotemporal dimensions, with $N$ being the number of frames, $w$ and $h$ being the width and height of the feature maps, and $c$ represents the feature dimension. This matrix encapsulates the spatiotemporal features extracted from the entire video sequence.
@@
We construct affinity matrix ${\bm{A}}=\mathbf{X}_{\text{FE}}\in\mathbb{R}^{d\times d}$ through feature entanglement, where $\mathbf{A}_{ij}$ represents the edge weight between nodes $i$ and $j$. This matrix captures the similarity between different nodes. The degree matrix $\mathbf{D}\in\mathbb{R}^{d\times d}$ is a diagonal matrix where $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$ represents the sum of edge weights connected to each node. The Graph Laplacian matrix $\mathbf{L}=\mathbf{D}-\mathbf{A}$ captures the topological structure of the graph and the differences between nodes.
@@
$\mathbf{A}^{(i)}_{jk}=\begin{cases}\mathbf{X}_{\text{FE}}^{(i)_{jk}},&\text{if}\mathbf{X}_{\text{FE}}^{(i)_{jk}}>q\times\overline{\mathbf{X}_{\text{FE}}^{(i)}}\\ 0,&\text{otherwise}\end{cases}$ (3)
@@
The resulting sparse adjacency matrix $\mathbf{A}^{(i)}$, along with the node features $\mathbf{X}^{(i)}$, serve as the input to the GCN for learning the graph structure and node representations. By combining the FE matrix ${\bm{X}}_{\text{FE}}$ and the sparse adjacency matrix ${\bm{A}}$, our method leverages the advantages of both representations, capturing rich spatiotemporal correlations while focusing on the most informative connections for DeepFake detection.
Graph Convolutional Networks (GCNs) have shown remarkable performance in various tasks by leveraging the power of graph-structured data. The core operation of GCNs in the $l$-th layer can be described as follows:
@@
where $\hat{\mathbf{A}}=\mathbf{A}+\mathbf{I}_{d}$ is the adjacency matrix with self-loops, $\hat{\mathbf{D}}_{ii}=\sum_{j=1}^{d}\hat{\mathbf{A}}_{ij}$ is the corresponding degree matrix, $\mathbf{W}^{(l)}$ is the weight matrix of the $l$-th layer, and $\sigma$ is the activation function. This equation describes the aggregation and transformation of node features based on the graph structure.
However, the performance of GCNs may be compromised when dealing with highly noisy scenarios, limiting their applicability in real-world situations. To address this challenge, we propose the incorporation of Graph Laplacian regularization into GCNs to enhance their robustness and improve their performance in the presence of significant noise.
Given an undirected graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges, the Graph Laplacian matrix $\mathbf{L}$ is defined as:
@@
where $\mathbf{D}$ is the degree matrix and $\mathbf{A}$ is the adjacency matrix of the graph $G$. The degree matrix $\mathbf{D}$ is a diagonal matrix, where $\mathbf{D}_{ii}$ equals the degree of node $i$ in $G$. It is worth noting that the Graph Laplacian matrix $\mathbf{L}$ is different from the matrix $\hat{\mathbf{L}}$ used in the GCN propagation rule, which will be discussed later.
As a real symmetric matrix, the Graph Laplacian matrix $\mathbf{L}$ possesses an eigendecomposition:
@@
where $\mathbf{U}$ is the matrix of eigenvectors and $\Lambda$ is the diagonal matrix of eigenvalues. The eigenvalues of $\mathbf{L}$ represent the frequencies of the graph signals, with smaller eigenvalues corresponding to lower frequencies and larger eigenvalues corresponding to higher frequencies. In GSP, the Graph Laplacian matrix is often used as a low-pass filter to smooth signals defined on graphs, effectively suppressing high-frequency noise while preserving low-frequency information. Although we do not explicitly compute the eigendecomposition of $\mathbf{L}$ in our implementation, it is essential to understand that the Graph Laplacian matrix inherently encodes the spectral properties of the graph, which enables effective feature smoothing and noise suppression.
In practice, we calculate the Graph Laplacian matrix $\mathbf{L}$ using the degree matrix $\mathbf{D}$ and the adjacency matrix $\mathbf{A}$, which are obtained from the graph structure. This calculation does not involve the explicit computation of eigenvalues and eigenvectors. However, the resulting matrix $\mathbf{L}$ still possesses the spectral properties that enable effective feature smoothing and noise suppression.
To integrate the Graph Laplacian smoothing prior into the GCN propagation rule, we propose a modified version of the Graph Laplacian matrix that incorporates the graph’s structural information and enables effective feature smoothing and noise suppression. Our proposed Graph Laplacian matrix, denoted as $\hat{\mathbf{L}}$, is defined as follows:
@@
The matrix $\hat{\mathbf{L}}$ is a normalized version of the Graph Laplacian matrix, which captures the graph’s structure and enables the smoothing of node features. By incorporating self-loops into the adjacency matrix, we ensure that each node’s feature is considered during the smoothing process, enhancing the stability and expressiveness of the learned representations.
To integrate the Graph Laplacian smoothing prior into the GCN propagation rule, we modify the propagation equation as follows:
@@
where $\mathbf{Z}^{(l)}$ represents the node features at layer $l$, $\mathbf{W}^{(l)}$ is the learnable weight matrix, and $\sigma$ is the activation function.
The term $\hat{\mathbf{L}}\mathbf{Z}^{(l)}$ effectively applies the Graph Laplacian smoothing prior to the node features. By multiplying the node features with the normalized Graph Laplacian matrix $\hat{\mathbf{L}}$, we achieve a smoothing effect that takes into account the graph’s structure. This operation allows the model to leverage the connectivity information encoded in the graph to refine the node features and suppress high-frequency noise.
It is important to note that although we do not explicitly compute the eigendecomposition of $\hat{\mathbf{L}}$ in our implementation, the matrix $\hat{\mathbf{L}}$ inherently possesses the spectral properties of the Graph Laplacian. By using $\hat{\mathbf{L}}$ in the GCN propagation rule, we implicitly leverage these spectral properties to achieve effective feature smoothing and noise suppression.
The effectiveness of the Graph Laplacian matrix $\hat{\mathbf{L}}$ as a low-pass filter can be understood by examining its eigendecomposition:
$\hat{\mathbf{L}}=\mathbf{U}\Lambda\mathbf{U}^{T},$ (9)
where $\mathbf{U}$ is the matrix of eigenvectors and $\Lambda$ is the diagonal matrix of eigenvalues. The eigenvalues of $\hat{\mathbf{L}}$ represent the frequencies of the graph signals, with smaller eigenvalues corresponding to lower frequencies and larger eigenvalues corresponding to higher frequencies. By multiplying the node features with $\hat{\mathbf{L}}$, we essentially apply a low-pass filter to the graph signals, attenuating the high-frequency components while preserving the low-frequency information. This operation effectively suppresses noise and promotes the smoothness of the node features across the graph. In practice, we compute $\hat{\mathbf{L}}$ using the normalized adjacency matrix $\hat{\mathbf{A}}$ and the corresponding degree matrix $\hat{\mathbf{D}}$, as shown in Equation (6). This computation can be efficiently performed using sparse matrix operations, without the need for explicit eigendecomposition.
To provide a rigorous theoretical analysis of the effectiveness of our proposed Graph Laplacian regularization, we can examine the convergence and generalization properties of the method. Let $\mathbf{Z}^{*}$ denote the optimal node features that minimize the loss function $\mathcal{L}(\mathbf{Z})$. We can show that by incorporating the Graph Laplacian regularization term $\hat{\mathbf{L}}\mathbf{Z}^{(l)}$ into the GCN propagation rule, the learned node features $\mathbf{Z}^{(l)}$ converge to $\mathbf{Z}^{*}$ under mild assumptions on the graph structure and the loss function. Specifically, if the graph is connected and the loss function is convex and smooth, the iterative updates of the node features using Equation (8) will converge to the optimal solution $\mathbf{Z}^{*}$ (see Theorem 1). Furthermore, the generalization error of the learned node features can be bounded by the graph Laplacian regularization term, indicating that the proposed method effectively controls the model complexity and prevents overfitting to noisy or irrelevant features.
By incorporating the Graph Laplacian smoothing prior into the GCN, our method effectively addresses the challenges posed by noisy scenarios. The smoothing operation helps to mitigate the impact of noisy or irrelevant features, enhancing the robustness and generalization ability of the learned representations.
@@
where ${\bm{W}}_{\text{cls}}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{cls}}}$ denotes the weight of the FC, $n_{\text{cls}}$ stands for the number of classes.
During the training phase, the cross-entropy loss is employed to optimize the model parameters:
$\mathcal{L}=-\sum_{c=1}^{n_{\text{cls}}}\mathbf{\hat{Y}}_{c}\log(\mathbf{Y}_{c})+\alpha|{\bm{X}}|_{1}$ (12)
where $\mathbf{Y}$ is the one-hot encoded ground-truth label, and $\alpha$ stands for weight of the sparsity constraint, where $\alpha=1e^{-5}$ in this study for all experiments.
By incorporating Graph Laplacian regularization into the GCN, our proposed method effectively addresses the challenges posed by noisy face sequences in DeepFake video detection. The Graph Laplacian smoothing helps to filter out noisy nodes and enhances the robustness and stability of the model. The combination of feature entanglement, sparse regularization, and Graph Laplacian regularization enables our method to make the most of the available valid information, suppress the influence of noisy features, and improve the correlation between relevant features. As a result, our approach achieves high-performance and robust DeepFake video detection, even in the presence of low-quality and noisy face sequences commonly encountered in real-world scenarios.
In summary, our proposed GCN with Graph Laplacian regularization effectively leverages the spectral properties of the Graph Laplacian matrix to achieve feature smoothing and noise suppression, without explicitly computing the eigendecomposition. By integrating the Graph Laplacian smoothing prior into the GCN propagation rule, our method enhances the robustness and generalization ability of the learned representations, making it highly suitable for detecting DeepFake videos in challenging real-world scenarios with noisy and unreliable face sequences. The theoretical analysis of the convergence and generalization properties of our method further validates its effectiveness and provides a solid foundation for its application.
@@ 3.1 Experimental Configuration
The robustness validation of the proposed method is the core of our investigation, particularly when applied to noisy face sequences containing many invalid faces. To achieve this, the use of representative benchmark datasets is essential. Therefore, we selected three well-established benchmark datasets for performance evaluation: FF++ [1], Celeb-DFv2 [3], and the large-scale DFDC dataset [2]. The FF++ dataset [1] comprises four distinct classes of manipulation methods: 1) DeepFakes (DF), 2) Face2Face (F2F), 3) FaceSwap (FS), and 4) NeuralTextures (NT). For each class, a set of 1,000 original videos was used to generate 1,000 manipulated versions, resulting in a total of 1,000 authentic and 4,000 doctored videos. The Celeb-DF dataset [3] contains 590 original videos and 5,639 manipulated counterparts, generated using improved generative adversarial networks at a resolution of $256\times 256$. To enhance the quality of the manipulated videos, a Kalman filter is employed in Celeb-DF [3] to mitigate temporal inconsistencies between successive frames. In addition to FF++ and Celeb-DF, we also utilize the DeepFake Detection Challenge (DFDC) dataset [2] to further validate the effectiveness of our proposed method. The DFDC dataset, created by Facebook in collaboration with other organizations, is a large-scale dataset designed to facilitate the development of DeepFake detection algorithms. It consists of over 100,000 videos, containing a mix of authentic and manipulated content generated using various state-of-the-art face swapping and facial reenactment techniques, ensuring a diverse and challenging set of DeepFakes for evaluation.
@@
Training Hyperparameters of Peer Methods. For performance evaluation, we compared our proposed method with several state-of-the-art DeepFake detection techniques, including MesoNet [5], Xception [6], $F^{3}$-net [32], RECCE [27], DFIL [28], UCF [30], CORE [26], and TALL-Swin [53]. The frame-level approaches, namely MesoNet, Xception, $F^{3}$-net, RECCE, DFIL, UCF, and CORE, were trained using the same strategy as described previously, with their default settings. However, the learning rates of Xception [6] and UCF [30] were adjusted to $2e^{-4}$ for better performance. The video-based approach, TALL-Swin [53], was trained using their default settings. During the training phase, we randomly selected $N$ facial images from the training set. The final authenticity verdict for the input video was determined by averaging the $N$ prediction outcomes corresponding to the $N$ facial images extracted from the input video, using a temporally centered cropping strategy. For all other methods, the number of frames $N$ used was set to $16$. The image size for Xception [6], $F^{3}$-net [32], RECCE [27], UCF [30], and CORE [26] is $256\times 256$, suggested by their default settings, while that for DFIL [28] and TALL-Swin [53] are $299\times 299$ and $224\times 224$, respectively.
@@
Furthermore, we assumed that each frame should contain at least one face to simulate adversarial attacks on face detectors in real-world scenarios. In cases where no face was detected in a frame, we replaced that frame with a black image, generating a noisy face sequence that allowed us to assess the robustness of GRACE under challenging conditions. Our experimental analysis employed three performance metrics: accuracy, Macro F1-Score, and Area Under the Receiver Operating Characteristic Curve (AUC). The Macro F1-Score accurately reflects the model’s performance under label imbalance situations. For simplicity, these metrics are referred to as Accuracy (Acc.), F1-Score, and AUC throughout the experimental sections.

### v3

@@ 3.3 GCN with Graph Laplacian Smoothing
GCNs have emerged as a powerful tool for processing graph-structured data [50, 51], making them a suitable choice for handling the graph embedding obtained through feature entanglement. However, the noise level in each node can vary depending on the degree of distortion. For highly distorted or noisy nodes (i.e., those with entangled features primarily contributed by invalid facial images), it is beneficial to eliminate them to ensure stable and excellent performance in DeepFake video detection. To mitigate the impact of features from invalid facial images, we judiciously integrate the GLSPR, a well-established concept in Graph Signal Processing (GSP) [52], as a regularizer into the GCN to filter out highly noisy nodes [53]. In this study, we ingeniously leverage the properties of GLSPR from GSP and seamlessly incorporate it into the GCN framework, enabling end-to-end training without explicitly computing the eigendecomposition of the Graph Laplacian matrix (GLm).
3.3.1 Graph Convolutional Network
@@
We construct affinity matrix $\mathbf{A}=\mathbf{X}_{\text{FE}}\in\mathbb{R}^{d\times d}$ through feature entanglement, where $\mathbf{A}_{ij}$ represents the edge weight between nodes $i$ and $j$. This matrix captures the similarity between different nodes. The degree matrix $\mathbf{D}\in\mathbb{R}^{d\times d}$ is a diagonal matrix where $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$ represents the sum of edge weights connected to each node.
@@
$\mathbf{A}^{(i)}_{jk}=\begin{cases}\mathbf{X}_{\text{FE},jk}^{(i)},&\text{if }\mathbf{X}_{\text{FE},jk}^{(i)}>q\times\overline{\mathbf{X}_{\text{FE}}^{(i)}}\\ 0,&\text{otherwise}\end{cases}$ (3)
@@
The resulting sparse adjacency matrix $\mathbf{A}^{(i)}$, along with the node features $\mathbf{X}^{(i)}$, serve as the input to the GCN for learning the graph structure and node representations. By combining the FE matrix $\mathbf{X}_{\text{FE}}$ and the sparse adjacency matrix $\mathbf{A}$, our method leverages the advantages of both representations, capturing rich spatiotemporal correlations while focusing on the most informative connections for the effective graph representation learning.
The core operation of GCN in the $l$-th layer can be described as follows:
@@
where $\hat{\mathbf{A}}=\mathbf{A}+\mathbf{I}_{d}$ is the adjacency matrix with self-loops, $\hat{\mathbf{D}}_{ii}=\sum_{j=1}^{d}\hat{\mathbf{A}}_{ij}$ is the corresponding degree matrix, $\mathbf{W}^{(l)}$ is the weight matrix of the $l$-th layer, $\mathbf{Z}^{(l)}$ represents the node features at layer $l$, and $\sigma$ is the activation function. This equation describes the aggregation and transformation of node features based on the graph structure.
3.3.2 Graph Laplacian Smoothing Prior Regularization
While GCNs demonstrate exceptional capability in modeling long-range dependencies through prior graph structures, their performance may be compromised when confronted with highly noisy scenarios, limiting their applicability in real-world situations. To address this challenge, we propose the incorporation of GLSPR into GCNs to enhance their robustness and improve their performance in the presence of significant noise. A noisy face sequence, potentially containing mis-detected faces, can result in inconsistent feature responses among neighboring nodes, thereby increasing the high-frequency components in the graph representation. GLSPR is particularly well-suited for attenuating these high-frequency components, effectively reducing the impact of noisy face sequences on DeepFake detection.
Given an undirected graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges, the GLm $\mathbf{L}$ is defined as:
@@ 3.3.2 Graph Laplacian Smoothing Prior Regularization
where $\mathbf{D}$ is the degree matrix and $\mathbf{A}$ is the adjacency matrix of the graph $G$. The degree matrix $\mathbf{D}$ is a diagonal matrix, where $\mathbf{D}_{ii}$ equals the degree of node $i$ in $G$.
As a real symmetric matrix, the effectiveness of the GLm $\hat{\mathbf{L}}$ as a low-pass filter can be understood by examining its eigendecomposition:
@@
where $\mathbf{U}$ is the matrix of eigenvectors and $\Lambda$ is the diagonal matrix of eigenvalues. The eigenvalues of $\mathbf{L}$ represent the frequencies of the graph signals, with smaller eigenvalues corresponding to lower frequencies and larger eigenvalues corresponding to higher frequencies. To elaborate, the high/low frequencies mentioned above differ from the conventional frequency-aware Deepfake detection methods [43]. In GSP [52], the GLm is often used as a low-pass filter to smooth signals defined on graphs, effectively suppressing high-frequency noise while preserving low-frequency information [53].
To integrate the GLSPR into the GCN propagation rule, we propose a modified version of the GLm that incorporates the graph’s structural information and enables effective feature smoothing and noise suppression, thereby reducing the temporal artifacts or semantic inconsistency on the spatial domain resulting from invalid facial images. Our proposed GLm, denoted as $\hat{\mathbf{L}}$, is defined as follows:
@@
The matrix $\hat{\mathbf{L}}$ is a normalized version of $\mathbf{L}$, which captures the graph’s structure and enables the smoothing of node features. By incorporating self-loops into the adjacency matrix, we ensure that each node’s feature is considered during the smoothing process, enhancing the stability and expressiveness of the learned representations. Afterwards, the propagation rule in Equation [4] of GCN can be modified as follows:
@@
The term $\hat{\mathbf{L}}\mathbf{Z}^{(l)}$ effectively applies the GLSPR to the node features. By multiplying the node features with the normalized GLm $\hat{\mathbf{L}}$, we achieve a smoothing effect that takes into account the graph’s structure. This operation allows the model to leverage the connectivity information encoded in the graph to refine the node features and suppress high-frequency noise.
It is important to note that although we do not explicitly compute the eigendecomposition of $\hat{\mathbf{L}}$ in our implementation, the matrix $\hat{\mathbf{L}}$ inherently possesses the spectral properties of the Graph Laplacian.
To examine the effectiveness of our proposed GLSPR, we provide a theoretical convergence analysis under simplified conditions in the Appendix and conduct the ablation study and empirical analyses (see Sec. 4.3). We can also examine the convergence and generalization properties of GLSPR. Let $\mathbf{Z}^{*}$ denote the optimal node features that minimize the loss function $\mathcal{L}(\mathbf{Z})$. We can show that by incorporating the GLSPR term $\hat{\mathbf{L}}\mathbf{Z}^{(l)}$ into the GCN propagation rule, the learned node features $\mathbf{Z}^{(l)}$ converge to $\mathbf{Z}^{*}$ under mild assumptions on the graph structure and the loss function. Specifically, if the graph is connected and the loss function is convex and smooth, the iterative updates of the node features using Equation (8) will converge to the optimal solution $\mathbf{Z}^{*}$ (see the Appendix).
@@
where ${\bm{W}}_{\text{cls}}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{cls}}}$ denotes the weight of the FC, $n_{\text{cls}}$ stands for the number of classes, $\hat{\mathbf{Y}}$ is the predicted result of the proposed GRACE for robust DeepFake video detection.
3.4 Sparsity Constraint on Network Optimization
While the affinity matrix $\mathbf{X}_{\text{FE}}$ effectively captures spatiotemporal dependencies within the feature representation, learning from noisy face sequences remains challenging. To enhance the network’s robustness, we introduce the SC as a regularization term in the learning objective, formulated as the $\ell_{1}$ norm of $\mathbf{X}$, denoted by $|\mathbf{X}|_{1}$. This constraint encourages the backbone to extract the relatively sparse feature representations from each input frame, thereby reducing the impact of noisy nodes, as illustrated in Figure 2. Consequently, the model focuses on the most informative features, improving its robustness to noisy face sequences and enhancing its efficacy in DeepFake video detection.
During the training phase, we optimize the model parameters using cross-entropy loss augmented with the SC:
$\mathcal{L}=-\sum_{c=1}^{n_{\text{cls}}}\mathbf{\hat{Y}}_{c}\log(\mathbf{Y}_{c})+\alpha|\mathbf{X}|_{1}$ (11)
where $\mathbf{Y}$ is the ground truth label, and $\alpha$ denotes the weight of the SC, set to $1e^{-5}$ across all experiments in this study.
The synergistic application of GLSPR and SC in our GRACE method yields robust performance in noisy sequences. By incorporating GLSPR into the GCN propagation rule, we enhance the robustness and generalization capability of the learned representations, making our approach particularly suitable for DeepFake video detection in challenging real-world scenarios with noisy and unreliable face sequences. Concurrently, SC guides the network to focus on essential information within $\mathbf{X}_{\text{FE}}$ and suppress irrelevant nodes, while enforcing the learning of sparse features for valid faces.
@@ 4.1 Experimental Configuration
The robustness validation of the proposed method is the core of our investigation, particularly when applied to noisy face sequences containing many invalid faces. To achieve this, the use of representative benchmark datasets is essential. Therefore, we selected three well-established benchmark datasets for performance evaluation: FF++ [1], Celeb-DFv2 [3], and the large-scale DFDC dataset [2]. The FF++ dataset [1] comprises four distinct classes of manipulation methods: 1) DeepFakes (DF), 2) Face2Face (F2F), 3) FaceSwap (FS), and 4) NeuralTextures (NT). For each class, a set of 1,000 original videos was used to generate 1,000 manipulated versions, resulting in a total of 1,000 authentic and 4,000 doctored videos. The Celeb-DF dataset [3] contains 590 original videos and 5,639 manipulated counterparts, generated using improved generative adversarial networks at a resolution of $256\times 256$. To enhance the quality of the manipulated videos, a Kalman filter is employed in Celeb-DF [3] to mitigate temporal inconsistencies between successive frames. The DFDC dataset [2], created by Facebook in collaboration with other organizations, is a large-scale dataset designed to facilitate the development of DeepFake detection algorithms. It consists of over 100,000 videos, containing a mix of authentic and manipulated content generated using various state-of-the-art face swapping and facial reenactment techniques, ensuring a diverse and challenging set of DeepFakes for evaluation.
@@
Training Hyperparameters of Peer Methods. For performance evaluation, we compared our proposed method with several state-of-the-art DeepFake detection techniques, including Xception [6], $F^{3}$-net [43], RECCE [38], DFIL [39], UCF [41], CORE [37], and TALL-Swin [40]. The image-based approaches, namely Xception, $F^{3}$-net, RECCE, DFIL, UCF, and CORE, were trained using the same strategy as described previously, with their default settings. However, the learning rates of Xception [6] and UCF [41] were adjusted to $2e^{-4}$ for better performance. The video-based approach, TALL-Swin [40], was trained using their default settings. During the training phase, we randomly selected $N$ facial images from the training set. The final authenticity verdict for the input video was determined by averaging the $N$ prediction outcomes corresponding to the $N$ facial images extracted from the input video, using a temporally centered cropping strategy. For all other methods, the number of frames $N$ used was set to $16$. The image size for Xception [6], $F^{3}$-net [43], RECCE [38], UCF [41], and CORE [37] is $256\times 256$, suggested by their default settings, while that for DFIL [39] and TALL-Swin [40] are $299\times 299$ and $224\times 224$, respectively.
@@
Furthermore, we assumed that each frame should contain at least one face to simulate adversarial attacks on face detectors in real-world scenarios. In cases where no face was detected in a frame, we replaced that frame with a black image, generating a noisy face sequence that allowed us to assess the robustness of GRACE under challenging conditions. Our experimental analysis employed three performance metrics: accuracy, macro F1-Score, and Area Under the Receiver Operating Characteristic Curve (AUC). For simplicity, these metrics are referred to as Accuracy (Acc.), F1-Score, and AUC throughout the experimental sections.